# What’s the Weight?

### v1

@@ 4.2 Results
Results from our simulation studies are given in Figure 2c-d and Supplemental Table LABEL:tab:simulation_results. Here, we discuss these results in the context of each simulation setting, where the data are generated to vary what sources of bias we expect to encounter when estimating the ACD.
4.2.1 Settings 1 and 3: No Bias Expected
In Settings 1 and 3, there is neither confounding nor selection bias in the data generating mechanism. Setting 1 provides a fair comparison across all the methods considered, as both the group propensity, $\Pr(A\mid X)$, and sample selection probabilities, $\Pr(S\mid A,X)$, were generated independent of any covariates. As shown, all of the methods except for the naive g-computation approach perform comparably in this setting with respect to the percent bias of the ACD estimator. Specifically, the percent bias among the three proposed estimators ranged from 0.499-0.647%, as compared to 0.166-0.538% among the existing methods (excluding the naive g-computation), suggesting that all of the methods considered accurately estimate the average controlled difference. This is expected, as the methods which do not account for confounding or selection bias are correctly specified, and the methods that do are unencumbered by the additional weights. The proposed approaches also had among the lowest MSEs, ranging from 0.475-0.480, as compared to 0.482-1.103 among the existing approaches. Lastly, across 200 independent replications, the proposed methods had coverage rates of 92.0-92.5%, as compared to 87.5-93.0% for the existing methods. Similarly, in Setting 3, where selection depends on $A$ but neither selection nor $A$ depend on $X$, all methods tend to perform well. Here, we see that the percent bias for the proposed methods ranged from 0.217-1.497%, on par with the other methods (0.138-0.417%). Mean squared errors ranged from 0.219-0.243 for the proposed methods, similar to MSEs of 0.222-0.547 among the existing methods. Further, all of the methods had nominal coverage rates of 94.5-99.0%.
4.2.2 Setting 2 and 4: Confounding Bias Expected
@@
In Settings 2 and 4, we generate $A$ conditional on $X$, with selection further being dependent on $A$ in Setting 4. In these settings, we now expect there to be confounding bias, as $\Pr(A\mid X)$ is no longer uniform over the range of $X$ values. As such, we see that all methods which account for both $A$ and $X$ tend to perform comparatively well, whereas the simple linear regression (SLR) of $Y$ on $A$ performs poorly due to confounding bias in $X$. Specifically, we see that the percent bias in the SLR estimator has increased from 0.138-0.166% in Settings 1 and 3 to 78.249-79.480% in Settings 2 and 4, whereas the other estimators (including the proposed methods) still see biases ranging in magnitude from 0.977-2.778% in Setting 2 and 0.210-1.526% in Setting 4. We further see an increase in MSE for the SLR estimator, with values ranging from 75.408-75.638, while the MSE for the other existing methods (0.404-1.176) and the proposed estimators (0.362-1.220) remained low. Once again, the coverage probabilities for the proposed (92.5-94.5%) and other existing (88.5-94.0%) remained at the nominal level, whereas the SLR did not cover the true ACD (0%).
4.2.3 Settings 5 and 7: Selection Bias Expected
@@
In Settings 5 and 7, we now generate $S$ conditional on $X$, with selection further being dependent on $A$ in Setting 7. Here, we now expect there to be selection bias induced by the $S\mid X$ relationship. In Setting 5, the percent bias among the existing approaches ranges from 9.182-9.544%, while the bias among the proposed methods ranges from 0.302-1.252%, demonstrating empirically that the proposed estimators are unbiased for the target ACD. Moreover, these results persist in Setting 7, with the percent bias among the existing approaches (3.992-6.809%) being higher than the proposed approaches (0.172-2.322%). Similar patterns can be seen when comparing the MSEs of the existing approaches (1.135-1.311 in Setting 5 and 0.587-0.663 in Setting 7) to those proposed in this work (0.236-0.325 in Setting 5 and 0.135-0.265 in Setting 7). Further, only the proposed methods achieved the nominal coverage rate (93.5-95.5% in Setting 5 and 91.0-97.0% in Setting 7), while the other methods show coverage ranging from 9.0-36.5% and 19.0-80.5% in Settings 5 and 7, respectively.
4.2.4 Settings 6 and 8: Both Confounding and Selection Bias Expected
@@
Lastly, Settings 6 and 8 represent the most complex scenarios of confounding and selection bias, with Setting 8 mirroring the conceptual diagram for our motivating data. In this setting, our proposed estimators outperform all current approaches with respect to the percent bias, MSE, and coverage for estimating the ACD. In particular, percent bias among existing approaches ranges from 7.847-82.816% in Setting 6 and 6.226-73.914% in Setting 8, whereas the bias among the proposed estimators ranges from 0.633-1.514% and 1.083-1.115%, respectively. This is expected as, the proposed methods have the correct balancing and generalizability weights under these settings, whereas the existing approaches fail to incorporate one or both weights, or fail to incorporate them correctly into the form of the estimator. The proposed approaches again had the lowest MSEs, ranging from 0.131-0.540 across both settings, as compared to 0.927-82.577 for the existing approaches. Lastly, the proposed approaches achieve nominal coverage (93.5-97.0% across both settings), whereas the existing methods do not (0-83.5%). For full results, see Supplemental Table LABEL:tab:simulation_results.
4.3 Sensitivity Analysis
In a sensitivity analysis, we study our proposed approaches under increasing model misspecification by introducing effect heterogeneity. Specifically, we systematically included larger interactions of race and covariates, where we varied $\gamma_{AX}\in\{0,0.01,0.05,0.1,0.5\}$. We estimate IPW1 and IPW2 as is, and we fit two versions of OM where we (1) correctly specify versus (2) misspecify the outcome model by not including the interaction term present in the data generating mechanism. Here we seek to assess the robustness of these approaches against the oracle estimator. The results from this analysis show that overall, all methods tend to perform well in terms of the bias, MSE, and coverage of the estimator, with OM performing acceptably under moderate misspecification. In particular, for $\gamma_{AX}\in\{0,0.01,0.05,0.1\}$, the method with the misspecified outcome model attained nominal coverage (ranging from 93.5-96.5%), with biases below 0.05% (0.008-0.028) and comparable MSEs to the other methods (0.001-0.003). However this approach incurs bias as treatment heterogeneity becomes more evident. Specifically, for the setting in which $\gamma_{AX}=0.5$, we see that the bias increases to 0.14% and the method does not cover the true ACD. These results suggest that the IPTW-type estimators (IPW1 and IPW2), which do not assume additivity, may be more robust. Full simulation results are presented in Supplemental Table LABEL:tab:sim_sensitivity.

### v2

@@ 4.2 Results
Results are given in Table LABEL:tab:simulation_results. Here, we discuss these results in the context of each simulation setting, where the data are generated to vary what sources of bias we expect to encounter when estimating the ACD.
Settings 1 and 3: No Bias Expected. When there is neither confounding nor selection bias in the data generating mechanism, most estimators recover the true ACD with negligible bias and nominal 95% coverage. Specifically, the relative bias among our four proposed estimators ranges from -0.002 to -0.001, as compared to -0.003 to 0.005% among the existing methods, as expected. The proposed approaches also have the lowest MSEs (0.012), as compared to MSEs ranging from 0.012-0.069 for the existing approaches, as well as standard errors that are comparable to the more efficient approaches (ASE: 0.109-0.113). Simple linear regression (SLR) has larger standard errors, though the ASE is within 5% of the MCSE, whereas the ASE of the traditional IPTW estimator is less than 50% that of its MCSE, suggesting the variance is underestimated. Lastly, the proposed methods have coverage rates of 95.6-96.2%, as compared to 95.6-96.4% for the existing methods except the IPTW estimator (58.6%). In contrast, the traditional IPTW estimator, though, unbiased on average, is particularly susceptible to underestimating the true variance when ignoring the stratified, clustered sampling design. in Setting 3, where selection depends on $A$, all methods except IPTW perform well. The relative biases for the proposed methods are all -0.005%, equal to the other methods, except IPTW (1.228). Mean squared errors are low (0.008) for the proposed methods, similar to MSEs of 0.008 to 0.024 among the existing methods, except IPTW (1.659). Further, all the methods had nominal coverage (95.2-96.4%), except IPTW (0%).
@@
Settings 2 and 4: Confounding Bias Expected. When we generate $A$ conditional on $X_{1}$ and $X_{2}$, with selection also depending on $A$ in Setting 4, we now expect there to be confounding bias. We see that methods which account for $A$ and $\boldsymbol{X}$ perform comparatively well, whereas SLR performs poorly, as $\Pr(A\mid X_{1},X_{2})$ is no longer uniform over the range of $\boldsymbol{X}$ values. Specifically, the relative bias in the SLR estimator increases from 0.002-0.005 in Settings 1 and 3 to 2.021-2.123 in Settings 2 and 4, whereas the other estimators (including the proposed methods) have biases ranging in magnitude from -0.004 to 0 in Setting 2 and 0.001 to 0.027 in Setting 4. We further see a higher in MSE for the SLR (1.089-1.357) and IPTW (0.078-1.421) estimators, while the MSE for the other existing methods (0.010-0.015) and the proposed estimators (0.010-0.022) remain low. Coverage for the proposed (93.4-94.4%) and other existing (91.4-93.6%) are also at the nominal level, whereas SLR (0%) and IPTW (0-63.4%) do not cover the true ACD. Among our proposed estimators, outcome modeling/direct standardization and AIPW deliver unbiased estimates with slightly lower ASE/MCSE than the purely IPW-based estimators, indicating better efficiency in this confounding scenario.
@@
Settings 5 and 7: Selection Bias Expected. When the $\boldsymbol{X}$ variables influence the probability of inclusion but do not confound the relationship between $Y$ and $A$, we expect there to be selection bias. In Setting 5, the relative bias among the existing approaches ranges from 0.001 to 0.231, with unweighted methods having higher bias (0.217-0.231) than the survey-weighted approaches (0.001-0.008). The bias among the proposed methods ranges from 0.000 to 0.027, demonstrating empirically that accounting for the survey weights corrects this selection bias when estimating the target ACD. These results persist in Setting 7, where selection also depends on $A$, with the relative bias among the existing unweighted approaches (0.173-3.560) being higher than existing survey-weighted approaches (0.001-0.005) and the proposed approaches (0.003-0.029). Similar patterns are seen in the MSEs of the existing approaches (0.019-0.162 in Setting 5 and 0.008-0.111 in Setting 7) versus the proposed methods (0.043-0.128 in Setting 5 and 0.018-12.637 in Setting 7). However, only the proposed methods have nominal coverage (94.2-97.4% in Setting 5 and 93.4-97.8% in Setting 7), while all other methods have coverage ranging from 19.4-85.8% and 0.0-90.2% in Settings 5 and 7, respectively. Among our proposed estimators, Outcome modeling again shows highest efficiency (ASE = 0.154 vs. MCSE = 0.137), as it is correctly specified, but AIPW approach strikes a balance between bias control and efficiency.
@@
Settings 6 and 8: Both Confounding and Selection Bias Expected. Lastly, Settings 6 and 8 are the most complex, with Setting 8 mirroring the conceptual diagram for our motivating data. In these settings, our proposed estimators outperform all current approaches with respect to the relative bias, MSE, and coverage for estimating the ACD. In particular, the relative biases for existing approaches that not use survey weights range from 0.203-1.198 in Setting 6 and 0.230-2.872 in Setting 8, and their coverages range from 0-44.6% and 0-14.4%, respectively. Further, while survey-weighting reduces bias (-0.004 to 0.021), the existing methods still under-cover (87.4-92.0%), indicating that design adjustments alone are insufficient when confounding and selection bias co-occur. In contrast, the bias among the proposed estimators ranges from -0.005 to 0.025, with coverages ranging from 94.6-98.2%. This is expected, as the proposed methods have the correct balancing and generalizability weights under these settings, whereas the existing approaches fail to incorporate one or both weights, or fail to incorporate them correctly into the form of the estimator. Among our proposed approaches outcome modeling/direct standardization is most efficient, while AIPW again achieves low bias, correct SEs, and nominal coverage, while offering better efficiency than the IPW-based approaches.
Table 2: Simulation results comparing different analytic methods for calculating the population average controlled difference (ACD) in $Y$ across AA for different data generating mechanism settings. Results presented are the true ACD (ACD) estimated ACD (Est. ACD), relative bias (Rel. Bias), average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE) and coverage probability (Cov.) averaged over 500 samples from the population generated under each setting.
Method ACD Est. ACD Rel. Bias ASE MCSE MSE Cov.
Setting 1: No Bias Expected ($\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=0$)
Oracle Estimator 1.003 1.000 -0.003 0.109 0.107 0.012 0.964
Simple Reg. 1.003 1.004 0.002 0.204 0.206 0.042 0.948
Multiple Reg. 1.003 1.001 -0.002 0.109 0.107 0.012 0.958
IPTW Estimator 1.003 1.008 0.005 0.110 0.263 0.069 0.586
Survey-Weighted Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.956
IPTW Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.958
IPTW + Survey-Weighted Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.956
Weighted IPTW + Survey-Weighted Multiple Reg. 1.003 1.001 -0.002 0.109 0.108 0.012 0.956
Outcome Modeling and Direct Standardization 1.003 1.001 -0.002 0.109 0.107 0.012 0.962
Inverse Probability Weighting 1 1.003 1.001 -0.002 0.110 0.108 0.012 0.956
Inverse Probability Weighting 2 1.003 1.002 -0.001 0.113 0.110 0.012 0.960
Augmented Inverse Probability Weighting 1.003 1.001 -0.002 0.110 0.108 0.012 0.958
Setting 2: Confounding Bias Expected ($\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=0$)
Oracle Estimator 0.996 0.995 0.000 0.115 0.120 0.014 0.930
Simple Reg. 0.996 2.021 1.030 0.195 0.194 1.089 0.000
Multiple Reg. 0.996 0.995 -0.001 0.115 0.121 0.015 0.932
IPTW Estimator 0.996 0.982 -0.014 0.135 0.280 0.078 0.634
Survey-Weighted Multiple Reg. 0.996 0.995 -0.001 0.116 0.122 0.015 0.932
IPTW Multiple Reg. 0.996 0.996 0.000 0.116 0.123 0.015 0.936
IPTW + Survey-Weighted Multiple Reg. 0.996 0.995 -0.001 0.117 0.124 0.015 0.928
Weighted IPTW + Survey-Weighted Multiple Reg. 0.996 0.995 0.000 0.117 0.124 0.015 0.928
Outcome Modeling and Direct Standardization 0.996 0.995 -0.001 0.116 0.121 0.015 0.934
Inverse Probability Weighting 1 0.996 0.993 -0.003 0.136 0.146 0.021 0.942
Inverse Probability Weighting 2 0.996 0.992 -0.004 0.138 0.147 0.022 0.940
Augmented Inverse Probability Weighting 0.996 0.996 0.000 0.118 0.124 0.015 0.934
Setting 3: No Bias Expected ($\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=0$)
Oracle Estimator 0.991 0.986 -0.005 0.090 0.090 0.008 0.962
Simple Reg. 0.991 0.996 0.005 0.169 0.167 0.028 0.962
Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.964
IPTW Estimator 0.991 2.267 1.288 0.105 0.177 1.659 0.000
Survey-Weighted Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.954
IPTW Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.966
IPTW + Survey-Weighted Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.952
Weighted IPTW + Survey-Weighted Multiple Reg. 0.991 0.986 -0.005 0.090 0.089 0.008 0.952
Outcome Modeling and Direct Standardization 0.991 0.986 -0.005 0.093 0.089 0.008 0.962
Inverse Probability Weighting 1 0.991 0.986 -0.005 0.093 0.089 0.008 0.962
Inverse Probability Weighting 2 0.991 0.986 -0.005 0.095 0.090 0.008 0.954
Augmented Inverse Probability Weighting 0.991 0.986 -0.005 0.093 0.089 0.008 0.960
Setting 4: Confounding Bias Expected ($\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$)
Oracle Estimator 0.969 0.969 0.001 0.094 0.100 0.010 0.938
Simple Reg. 0.969 2.123 1.191 0.167 0.161 1.357 0.000
Multiple Reg. 0.969 0.995 0.027 0.097 0.103 0.011 0.930
IPTW Estimator 0.969 2.146 1.215 0.148 0.188 1.421 0.000
Survey-Weighted Multiple Reg. 0.969 0.970 0.001 0.094 0.100 0.010 0.940
IPTW Multiple Reg. 0.969 0.995 0.027 0.098 0.106 0.012 0.914
IPTW + Survey-Weighted Multiple Reg. 0.969 0.995 0.027 0.098 0.107 0.012 0.920
Weighted IPTW + Survey-Weighted Multiple Reg. 0.969 0.969 0.001 0.095 0.102 0.010 0.934
Outcome Modeling and Direct Standardization 0.969 0.970 0.001 0.096 0.100 0.010 0.942
Inverse Probability Weighting 1 0.969 0.980 0.011 0.111 0.112 0.013 0.944
Inverse Probability Weighting 2 0.969 0.976 0.008 0.118 0.123 0.015 0.934
Augmented Inverse Probability Weighting 0.969 0.970 0.001 0.098 0.102 0.010 0.942
Setting 5: Selection Bias Expected ($\tau_{X}=0$, $\beta_{A}=0$, $\beta_{X}=1$)
Oracle Estimator 0.971 0.976 0.006 0.133 0.137 0.019 0.946
Simple Reg. 0.971 1.187 0.222 0.136 0.127 0.063 0.650
Multiple Reg. 0.971 1.182 0.217 0.074 0.074 0.050 0.194
IPTW Estimator 0.971 1.195 0.231 0.075 0.279 0.128 0.280
Survey-Weighted Multiple Reg. 0.971 0.973 0.002 0.163 0.214 0.046 0.858
IPTW Multiple Reg. 0.971 1.182 0.218 0.074 0.074 0.050 0.194
IPTW + Survey-Weighted Multiple Reg. 0.971 0.972 0.001 0.163 0.216 0.046 0.866
Weighted IPTW + Survey-Weighted Multiple Reg. 0.971 0.978 0.008 0.159 0.207 0.043 0.858
Outcome Modeling and Direct Standardization 0.971 0.979 0.008 0.154 0.137 0.019 0.974
Inverse Probability Weighting 1 0.971 0.971 0.000 0.197 0.211 0.044 0.942
Inverse Probability Weighting 2 0.971 0.944 -0.027 0.327 0.402 0.162 0.954
Augmented Inverse Probability Weighting 0.971 0.971 0.000 0.209 0.227 0.051 0.946
Setting 6: Both Confounding and Selection Bias Expected ($\tau_{X}=1$, $\beta_{A}=0$, $\beta_{X}=1$)
Oracle Estimator 1.002 1.020 0.018 0.116 0.118 0.014 0.946
Simple Reg. 1.002 2.201 1.198 0.128 0.121 1.454 0.000
Multiple Reg. 1.002 1.225 0.223 0.079 0.078 0.056 0.192
IPTW Estimator 1.002 1.205 0.203 0.133 0.336 0.154 0.446
Survey-Weighted Multiple Reg. 1.002 1.015 0.013 0.144 0.168 0.028 0.898
IPTW Multiple Reg. 1.002 1.230 0.228 0.080 0.081 0.059 0.202
IPTW + Survey-Weighted Multiple Reg. 1.002 1.018 0.016 0.147 0.187 0.035 0.874
Weighted IPTW + Survey-Weighted Multiple Reg. 1.002 1.023 0.021 0.143 0.176 0.031 0.868
Outcome Modeling and Direct Standardization 1.002 1.021 0.019 0.136 0.119 0.015 0.978
Inverse Probability Weighting 1 1.002 1.003 0.002 0.187 0.183 0.033 0.956
Inverse Probability Weighting 2 1.002 1.026 0.025 0.261 0.307 0.095 0.982
Augmented Inverse Probability Weighting 1.002 1.022 0.020 0.199 0.244 0.060 0.960
Setting 7: Selection Bias Expected ($\tau_{X}=0$, $\beta_{A}=1$, $\beta_{X}=1$)
Oracle Estimator 0.997 0.991 -0.007 0.094 0.091 0.008 0.956
Simple Reg. 0.997 1.170 0.173 0.097 0.093 0.038 0.562
Multiple Reg. 0.997 1.205 0.208 0.053 0.051 0.046 0.030
IPTW Estimator 0.997 4.548 3.560 0.095 0.171 12.637 0.000
Survey-Weighted Multiple Reg. 0.997 1.000 0.002 0.122 0.137 0.019 0.902
IPTW Multiple Reg. 0.997 1.205 0.208 0.053 0.051 0.046 0.030
IPTW + Survey-Weighted Multiple Reg. 0.997 0.998 0.001 0.122 0.140 0.020 0.898
Weighted IPTW + Survey-Weighted Multiple Reg. 0.997 1.003 0.005 0.120 0.134 0.018 0.902
Outcome Modeling and Direct Standardization 0.997 0.992 -0.005 0.110 0.092 0.008 0.978
Inverse Probability Weighting 1 0.997 1.002 0.005 0.148 0.140 0.020 0.954
Inverse Probability Weighting 2 0.997 1.026 0.029 0.262 0.332 0.111 0.934
Augmented Inverse Probability Weighting 0.997 1.001 0.003 0.158 0.153 0.023 0.968
Setting 8: Both Confounding and Selection Bias Expected ($\tau_{X}=1$, $\beta_{A}=1$, $\beta_{X}=1$)
Oracle Estimator 1.001 0.991 -0.010 0.097 0.093 0.009 0.956
Simple Reg. 1.001 2.210 1.207 0.117 0.110 1.474 0.000
Multiple Reg. 1.001 1.232 0.230 0.073 0.079 0.059 0.126
IPTW Estimator 1.001 3.877 2.872 0.173 0.172 8.301 0.000
Survey-Weighted Multiple Reg. 1.001 0.994 -0.007 0.118 0.134 0.018 0.920
IPTW Multiple Reg. 1.001 1.233 0.231 0.074 0.081 0.060 0.144
IPTW + Survey-Weighted Multiple Reg. 1.001 1.016 0.015 0.109 0.123 0.015 0.912
Weighted IPTW + Survey-Weighted Multiple Reg. 1.001 0.998 -0.004 0.122 0.145 0.021 0.912
Outcome Modeling and Direct Standardization 1.001 0.993 -0.008 0.114 0.095 0.009 0.982
Inverse Probability Weighting 1 1.001 0.996 -0.005 0.157 0.155 0.024 0.954
Inverse Probability Weighting 2 1.001 1.017 0.016 0.226 0.264 0.070 0.968
Augmented Inverse Probability Weighting 1.001 0.992 -0.010 0.154 0.158 0.025 0.946
4.3 Sensitivity Analyses
We carry out five sensitivity analyses to better understand the robustness of our approach to the following various real-world situations. In the following, we briefly describe the setup of each sensitivity analysis and our key takeaways. Full sensitivity analysis results can be found in Appendix B.
4.3.1 Single-Stage Unequal Probability Sampling
We compare our estimators to the current standard methods, but with the data drawn from a single-stage unequal probability sample, to examine how the performances of the methods change under different sampling designs. In this sensitivity analysis, we set the number of strata and clusters to be $n_{strata}=n_{cluster}=1$ and retain all other parameter settings from our main simulation.
When the data are generated under a single-stage unequal probability sample, all estimators show reduced Monte Carlo variability compared to our main, stratified clustered complex design setting, but their relative performances remain the same. The unweighted and survey-weighted regression methods are unbiased across all eight scenarios, with analytic SEs matching MCSEs and coverage near 95% in the earlier settings, with performance decreasing under settings with confounding and/or selection bias. In contrast, the four proposed methods have low bias, within $\pm$ 1% of the truth, ASEs closely aligned with the MCSEs, and coverages of 94-98% across all settings. Outcome Modeling often achieves the lowest ASE and MCSE, while IPW1 and IPW2 remain robust with minimal bias and nominal coverage, and AIPW combines these results to offer reliable inference even when either the propensity or outcome model alone is correctly specified. Full results are given in Supplemental Table B.1.
4.3.2 Propensity/Outcome Model Misspecification
In a second sensitivity analysis, we study our proposed approaches under increasing misspecification of either/both the propensity and outcome models. We consider nine scenarios by crossing three levels of propensity score model misspecification (via the interaction effect, $\tau_{X12}\in\{0,0.1,0.5\}$) with three levels of outcome model misspecification (via the group heterogeneity effect, $\alpha_{AX}\in\{0,0.1,0.5\}$). In each case, we estimate only main effects models, thereby omitting the true interaction terms when they are non-zero. The results of this sensitivity analysis show that outcome modeling alone is vulnerable to both mild and moderate outcome model misspecification, as expected. Further, IPW1 is fairly robust to moderate propensity score misspecification but loses efficiency, while IPW2 is unstable under propensity score misspecification. The AIPW estimator remains the only estimator with low bias in most settings, reasonable variance, and coverage closest to 95% across the fully factorial design.
Specifically, when neither model is misspecified, all four proposed methods match the oracle’s performance (bias $\approx 0$, ASE $\approx$ MCSE, and coverage at the nominal level), with outcome modeling/direct standardization being most efficient, and AIPW and IPW1 being slightly more efficient than IPW2. Under mild outcome model misspecification, outcome modeling incurs 17% bias and has only 68% coverage, while IPW1, IPW2, and AIPW are unbiased and having nominal coverage. These results are magnified in the moderate outcome misspecification setting. When only the propensity model is mildly or moderately misspecified, outcome modeling and AIPW remain unbiased, and IPW1 is fairly robust, but IPW2 shows large bias under moderate misspecification (-20%) despite nominal coverage (94%). In scenarios where both models are mildly or mixed misspecified, AIPW consistently has bias near 2-3% and coverage $\geq$ 88%, IPW1 maintains < 3% bias and $\geq$ 91% coverage, whereas outcome modeling and IPW2 break down under moderate misspecification. Finally, when both models are moderately misspecified, AIPW’s bias is 7.9% (coverage 83%) and IPW1’s bias is 6.1% (coverage 85%), while outcome modeling and IPW2 have bias > 20%. Overall, the doubly robust approach offers the best balance of bias control and inference validity across all misspecification patterns, as expected. Full results are given in Supplemental Table B.2.
4.3.3 Limited Overlap
In practice, the positivity assumption may be violated due to limited overlap in propensity scores between groups. To study the impact of this violation, we conduct a sensitivity analysis exploring scenarios with different degrees of overlap. Following Zeng et al. (2025), we tune the parameters in the propensity model generating mechanism to simulate propensity scores with good versus poor overlap. Specifically, we vary $\tau_{A}\in\{0.5,2.0\}$, in the true propensity model to emulate settings with good versus poor overlap in the propensity scores, respectively. Figure 2 illustrates the distributions of the propensity scores under these scenarios.
Under good overlap ($\tau_{A}=0.5$), all four proposed methods yield unbiased ACD estimates and nominal coverage. Outcome modeling has the smallest bias (-0.2%) and ASE (0.113), while IPW1 and IPW2 have modestly higher bias (0.7% and 2.8%) and (0.153, 0.213). AIPW balances these outcomes, with 0.8% bias, ASE = 0.150, and 96.8% coverage. When overlap is poor ($\tau_{A}=2.0$), outcome modeling remains unbiased (1.4%), has low ASE (0.129) that aligns with its MCSE (0.109), and maintains nominal coverage (97.2%). Similarly, AIPW maintains low bias (-1.0%) and nominal coverage (96.2%). In contrast, IPW1 and IPW2 have higher bias (-9.5% and 2.5%), underestimated standard errors (ASE of 0.542 and 0.352 versus and MCSE of 0.966 and 0.714, respectively), reducing their coverage to 91.8% and 90.6%. These results suggest that while the IPW-based estimators are affected by poorer overlap, outcome modeling and AIPW are robust and provide more stable performance in terms of bias, variability, and coverage. For the full simulation results, see Supplemental Table B.3.
Figure 2: Distribution of estimated propensity scores by group membership under two overlap scenarios. A. Good Overlap $(\tau_{A}=0.5)$. Propensity scores for the comparison group (purple, $A=0$) and group of interest (teal, $A=1$) show good overlap around 0.5, supporting stable inverse-probability weighting. B. Poor Overlap $(\tau_{A}=2.0)$. Propensity scores concentrate near 0 for the comparison group and near 1 for the group of interest, indicating limited common support.
4.3.4 Estimated Selection Mechanism
As discussed in Section 3.4.2, if sample selection mechanism is not known, we need to model $\Pr(S\mid A,\boldsymbol{X})$. In this supplementary analysis, we treat the probability of selection given our variable of interest and our confounders as unknown and estimate $\Pr(S\mid A,\boldsymbol{X})$ via a beta regression, as recommended by Elliot (2009). Note that this deviates from our motivating example, but is relevant for applications with non-probability samples. Here, we replicate all the settings of our main simulation, but for our proposed methods we treat the selection probabilities as unknown and estimate them.
These results show that our proposed approaches still have low bias and again are the only methods to achieve nominal coverage in our setting of interest. Specifically, outcome modeling remains nearly unbiased (relative bias within $\pm$ 1%) with ASEs that align with the Monte Carlo SEs and 94-96% coverage. Both IPW1 and IPW2 also maintain minimal bias ($\leq$ 1%) and nominal coverage (92-96%), although their Monte Carlo variability is occasionally slightly larger than their ASEs in more complex settings. AIPW delivers the most consistent performance, combining low bias ($\pm$ 1%) with ASEs within 10% of the MCSEs and coverage around 95% in every scenario. Moreover, while the standard survey-weighted regressions also achieve low bias, they all have coverage $\leq$ 90.4% in our setting of interest. Full results are given in Supplemental Table B.4.
4.3.5 Sampling Covariate Omission
Lastly, we extend our main simulation by introducing two additional, unobserved covariates, $U_{1}$ and $U_{2}$, that vary by stratum and cluster and drive the selection mechanism, but do not enter the true propensity or outcome models. In generating each sample, we first simulate $U_{1}$ and $U_{2}$ with stratum- and cluster-specific means and variances, then let the probability of inclusion depend on $A$, $X_{1}$, $X_{2}$, $U_{1}$, and $U_{2}$, with coefficients $\beta_{U_{1}}=\beta_{U_{2}}=1$ across all settings. All other parameter settings followed from the main simulation. In our analysis, however, we treat the selection mechanism as unknown and estimate $\Pr(S=1\mid A,\boldsymbol{X})$, omitting $U_{1}$ and $U_{2}$ to understand if this induces bias. We then apply our four proposed estimators using the estimated selection model, and compare their performance to the oracle estimator and the existing methods.
Despite this misspecification, all four proposed methods retain low bias and nominal coverage in most settings. Outcome modeling incurs a modest negative bias (ranging from –3.3% to –5.7% across all scenarios), but it remains efficient (MCSE from 0.06 to 0.12) and maintains coverage at the nominal level. IPW1, which uses a single estimated selection model, shows slightly larger bias (-1.7% to -10.6%) with coverage generally between 86% and 96%. IPW2, which estimates more models, exhibits bias within $\pm$ 3.6% in most scenarios and preserves coverage around 95-97%, though its variance inflation is substantial in some settings. Lastly, AIPW strikes the best balance, with bias remaining under 8.1% (-8.1% to 6.7%), ASEs aligning closely with the MCSEs, and coverage between 90.8% and 95.4%. Overall, even when key design covariates are omitted from the estimated selection model, the proposed approaches have reasonable performance, with the AIPW estimator being particularly robust to this form of misspecification. Full results are given in Supplemental Table B.5.